callmeizzy1234:

need hlp remembering perfect squares 1-12

callmeizzy1234:

example give me a number 144 then i give you the square of that number so for 144 its 12

AZ:

what is 1 * 1 = 2 * 2 = 3 * 3 = 4 * 4 = 5 * 5 = and so on

AZ:

ooo you want us to quiz you?

callmeizzy1234:

yes pls

AZ:

What is the square of 64

callmeizzy1234:

8

AZ:

What about the square root of 25

callmeizzy1234:

5

AZ:

How about 42

callmeizzy1234:

idk

AZ:

Exactly! Because it's not a perfect square

callmeizzy1234:

okay phew

AZ:

What about 81

callmeizzy1234:

9

AZ:

and 16?

callmeizzy1234:

4

AZ:

and 49

callmeizzy1234:

7

AZ:

Good good!! And is 50 a perfect square?

callmeizzy1234:

nope

AZ:

Good!! What about 100?

callmeizzy1234:

10

AZ:

Perfect! And 4?

callmeizzy1234:

2

AZ:

How about 120

callmeizzy1234:

not a perfect square

callmeizzy1234:

but its close 121 is 11

AZ:

Yes!!! Good job!!

AZ:

what about 36

callmeizzy1234:

6

AZ:

And finally, what about 9

callmeizzy1234:

not a pefect square

AZ:

uh oh are you sure?

callmeizzy1234:

oh 3

AZ:

Now you got it!

callmeizzy1234:

🤦

callmeizzy1234:

and then 1 is 1

AZ:

yup!

callmeizzy1234:

so thats 1-12 right